OpenStudy (anonymous):

I am so confused... The parametric equations x= t^2, y=t^4 have the same graph as x=t^3, y=t^6 ?

hartnn (hartnn):

thats correct , if u observe , from x= t^2 and y= t^4 u have y= x^2 right?? similarly from x=t^3 and y = t^6 , u will have y= x^2 hence both will have same graph of y=x^2 makes sense??

OpenStudy (anonymous):

oooook thanks!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

in the first case the graph is always confined in the first quadrant but in the second case the graph is cofined to both the first and the second quadrant and is symmetric about upper half of y-axis...

hartnn (hartnn):

how did u conclude 'in the first case the graph is always confined in the first quadrant' @matricked ? ? i think it will be same as 2nd graph u mentioned.

OpenStudy (anonymous):

whatever real values t takes x and y are both positive in the first case but if t is negative in the second case x is negative though it is that for both cases y= x^2